OpenStudy (anonymous):

Solve y = -2x + 3 for x

OpenStudy (anonymous):

y-3=-2x y-3/-2=x

OpenStudy (anonymous):

y=-2x+3 so y-3=-2x thus x=y-3/-2

OpenStudy (texaschic101):

y = -2x + 3 (subtract 3 from both sides) y - 3 = -2x (now divide both sides by -2) (y - 3)/-2 = x or -1/2y + 3/2 = x

OpenStudy (anonymous):

idk these are the answer choices x = y/-2 - 3 x = y + 3/ -2 x = y – 3/ -2 x = y/3 + 2

OpenStudy (anonymous):

please help

OpenStudy (anonymous):

The third answer is right.

OpenStudy (texaschic101):

x = (y - 3) / -2

OpenStudy (anonymous):

thank you so much @Chineseboy15 and @texaschic101 :)

OpenStudy (anonymous):

You are welcome.